OpenStudy (anonymous):

Solve the inequality. 3/5x<-11

OpenStudy (anonymous):



OpenStudy (anonymous):

i know that you have to multiply each side to it's reciprocal which is \[\frac{ 5 }{ 3 }\]

OpenStudy (anonymous):

The options are 3 2 -3 or none of the above

OpenStudy (anonymous):

then after that i'm stuck don't know how to do the multiplication

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

but how did you get those?

OpenStudy (jiteshmeghwal9):

\[\dfrac{3}{5}x \le -11\]multiply 5 both sides & then divide 3 both sides :)

OpenStudy (anonymous):

it gives me -18.3 :/

OpenStudy (anonymous):

or is it 5?

OpenStudy (jiteshmeghwal9):

\[x \le -\left( \dfrac{55}{3} \right)\]

OpenStudy (jiteshmeghwal9):

it is -18.3

OpenStudy (anonymous):

just one question why did you multiply 5 by each side and not 5/3?

OpenStudy (jiteshmeghwal9):

It's ur choice whether u multiply 3/5 both sides or only 5. Answer will be same :)